Find the number of years needed for a $3,470 deposit to grow to $4,611 when interest is 5.35% compounded daily. 


1
Expert's answer
2021-07-14T10:12:30-0400

First, convert R as a percent to r as a decimal

"r = \frac{5.35}{100}\\

r = 0.0535" rate per year,

"A = P(1 +\frac{ r}{n})^{nt}"

A = Accrued amount,P is the Principal ,r is annual nominal interest rate as a decimal,n is number of compounding periods per unit of time and t time in decimal years

"4611 = 3470\times(1 + \frac{0.0535}{365})^{(365)(t)}\\
4611 = 3470\times(1.000146575)^{(365t)}\\
\frac{4611}{3470}=(1.000146575)^{(365t)}\\
1.328818444=(1.000146575)^{(365t)}\\
t=\frac{ln(1.328818444)}{365ln(1.00014657)}\\t=5.31423659
=5.3years"